OpenStudy (shamil98):

Solve the initial-value problem: \[\huge 4y'' - 4y' + y = 0 \] \[\huge y(0) = 1 , y'(0) = -1.5\]

OpenStudy (shamil98):

The auxillary equation would be: \[4r^2 - 4r + 1 = 0\] The root is r = 1/2 \[\large y = c_1e^{1/2x} + c_2 xe^{1/2x}\]

OpenStudy (isaiah.feynman):

What a curious high school kid.

OpenStudy (anonymous):

Characteristic polynomial is \(\large 4r^2-4r+1 = 0\implies (2r-1)^2=0\implies r= \frac{1}{2}\) with multiplicity 2. Thus, \(\large y = c_1e^{x/2} + c_2xe^{x/2}\). Now, \(\large y^{\prime} = \frac{1}{2}c_1e^{x/2} + c_2(e^{x/2}+\frac{1}{2}xe^{x/2})\). Applying the initial conditions leaves us with \[\large \left\{\begin{aligned}c_1 &= 1\\ \frac{1}{2}c_1 + c_2 &= -\frac{3}{2}\end{aligned}\right.\implies c_1=1\quad\text{and}\quad c_2 = -2\]Therefore \(\large y=e^{x/2} -2xe^{x/2}\).

OpenStudy (shamil98):

\[y'(x) = (1/2)e^{1/2x} + c_2(e^{x/2} + 1/2xe^{x/2})\] \[1= c_1\] \[-1.5 = \frac{ 1 }{ 2 }c_1 e^{x/2} + c_2 (e^{x/2} + 1/2x e^{x/2})\]

OpenStudy (shamil98):

yeah, i was learning some calc 3 today lol

OpenStudy (shamil98):

I'm curious though, how did @ChristopherToni you evaluate c_2 so fast? :o

OpenStudy (anonymous):

plug in 0 into \(\large y^{\prime}\) and it should simplify to \(\large\frac{1}{2}c_1 + c_2 = -\frac{3}{2}\).

OpenStudy (anonymous):

and since \(\large c_1=1\) then we have \(\large \frac{1}{2} + c_2 = -\frac{3}{2} \implies c_2=-\frac{4}{2}=-2\).

OpenStudy (shamil98):

\[-1.5 = 1/2 c_1 + c_2 (1)\] \[-1.5 = 1/2 + c_2\] \[c_2 = -2\]

OpenStudy (shamil98):

yeah got it now xD, brain is slow, i need sleep lol

OpenStudy (shamil98):

\[y = c_1e^{1/2x}+c_2xe^{1/2x}\] pluggin em in c_ 1 = 1 c_2 = -2 \[y = e^{\frac{ 1 }{ 2 }x} - 2xe^{\frac{ 1 }{ 2 }x}\]

OpenStudy (anonymous):

Good job. :D

OpenStudy (shamil98):

ty, learned this topic today, just seeing if i still remember xD